OpenStudy (anonymous):

points (1,1) and (5,2) lie on line k. what is the slope of the line that is perpendicular to k?

OpenStudy (angela210793):

line k has \[slope=(y2-y1)/(x2-x1)=(2-1)/(5-1)=1/4\] the slope of the line perpendicular to k is: slope=-1/k=-1/(1/4)=-4